2. In an isolated housing area consisting of 50 households, the residents are allowed to have at most one dog and at most one cat per household. Currently 25 households have a dog, 7 have both a cat and a dog, and 18 have neither a cat nor a dog. An experiment consists of randomly selecting a household and observing whether that household has a cat or a dog, or both. Using D to denote the event that the selected household has a dog and C to denote the event that the selected household has a cat, symbolically denote the following events and identify the number of households represented by eacha The selected household has a dog but no cat. b The selected household has a cat. с .lhe: seleci:ed h()usehold has a cai: but no d()g. d The selected household has a cat or a dog but not both.
Answer #1

here from given information:

n(D) =25 (as 25 house have dogs) ; n(C n D)= 7 (as 7 have both cats and dogs)

n(C' n D') =18

as n(C u D) =50-n(C' n D')=50-18=32

a)

n(D n Cc )=n(D) -n(D n C)=25-7=18

b)

n(C) =n(C u D)+n(C n D)-P(D)=32+7-25=14

c)

n(C n D') =n(C)-n(C n D)=14-7=7

d) n(C u D)-n(C n D)=32-7=25